# Efficient Classical Algorithms for Simulating Gaussian Boson Sampling on Graphs

### v1

@@ Abstract
Gaussian Boson Sampling (GBS) is a promising candidate for demonstrating quantum computational advantage and can be applied to solving graph-related problems. In this work, we propose Markov chain Monte Carlo-based algorithms to simulate GBS on undirected, unweighted graphs. Our main contribution is a double-loop variant of Glauber dynamics, whose stationary distribution matches the GBS distribution. We further prove that it mixes in polynomial time for dense graphs using a refined canonical path argument. Numerically, we conduct experiments on graphs with 256 vertices, larger than the scales in former GBS experiments as well as classical simulations. In particular, we show that both the single-loop and double-loop Glauber dynamics improve the performance of original random search and simulated annealing algorithms for the max-Hafnian and densest $k$-subgraph problems up to 10x. Overall, our approach offers both theoretical guarantees and practical advantages for classical simulations of GBS on graphs.
preprint: APS/123-QED
@@ I Introduction
On the other hand, classical simulation algorithms for Gaussian Boson sampling have been actively explored. Oh et al. [13] developed a classical algorithm for boson sampling based on dynamic programming, which exploits the limited connectivity of the linear-optical circuit to improve efficiency. Separately, Oh et al. [14] proposed a tensor-network-based classical algorithm to simulate large-scale Gaussian Boson sampling experiments with photon loss, requiring relatively modest computational resources. In the context of graph-theoretical applications of GBS, an important property is that the adjacency matrix encoded in the Gaussian boson sampler is non-negative, which is believed to make the problem more tractable than in the general case. Oh et al. [15] took advantage of this property to design a quantum-inspired classical algorithm for finding dense subgraphs and their numerical results suggest that the advantage offered by a Gaussian boson sampler is not significant. However, an open question remains whether a classical algorithm for GBS on general graphs with provable performance guarantees can achieve a computational cost comparable to that of a Gaussian boson sampler.
@@
In this work, we adopt Markov chain Monte Carlo (MCMC) algorithms to simulate Gaussian Boson sampling on graphs. MCMC is a standard class of sampling algorithms with well-established theoretical guarantees [16, 17]. Among MCMC methods, Glauber dynamics [18] is particularly widespread due to its simplicity and rigorous analytical foundations. Glauber dynamics generates samples from the matchings of an undirected and unweighted graph by iteratively adding or removing edges with biased transition probabilities. The resulting stationary distribution is proportional to the power of the number of edges in the matchings. For each matching sampled from Glauber dynamics, we consider its support vertex set as a sampled subset of vertices. The probability of sampling a given vertex set is further weighted by the number of perfect matchings within that set, which is also the Hafnian of the adjacency matrix of the graph.
@@
We propose a double-loop Glauber dynamics with a rigorous theoretical guarantee that its stationary distribution is identical to the sampling distribution of GBS on graphs. Specifically, unlike the standard single-loop Glauber dynamics, which yields a stationary distribution proportional to the Hafnian of subgraphs, the double-loop approach ensures a stationary distribution proportional to the square of the Hafnian, which coincides with the sampling distribution of GBS. Concretely, in the double-loop Glauber dynamics, when deciding whether to remove an edge, we run a secondary Markov chain to uniformly sample a perfect matching from the current subgraph. The removal decision is then based on whether this edge reappears in the newly sampled matching. Furthermore, for dense graphs, we prove that the double-loop Glauber dynamics has a polynomial mixing time, demonstrating its computational feasibility in these cases. The key to demonstrate the rapid mixing lies in the canonical path technique introduced in [16], which routed flows between every pair of matchings without creating particularly congested “pipes”, and bounded the mixing time by estimating the maximum congestion of possible transitions. More specifically, for complete graphs, we establish an enhanced canonical path framework that permits multiple paths between any two matchings. By constructing specially designed paths with favorable symmetry properties to distribute flows efficiently, we ultimately estimate maximum congestion through calculating total congestion of symmetrical transitions. For dense graphs, the maximum congestion is bounded by estimating the ratio between the congestion of dense graphs and complete graphs, thereby yielding a polynomial mixing time.
Our numerical simulations confirm that both single-loop and double-loop Glauber dynamics improve the performance of original random search and simulated annealing algorithms for the max-Hafnian and densest $k$-subgraph problems, providing empirical validations of our theoretical findings. Our experiments are conducted on graphs with 256 vertices, larger than the scales in former GBS experiments [12] as well as classical simulations [14]. In verification, the variants enhanced by Glauber dynamics are up to 3x better than the original classical algorithms. On random graphs, the enhanced variants achieve score advantage up to 4x. On bipartite graphs, the enhanced variants are up to 10x better than the original classical algorithms.
@@
The rest of the paper is organized as follows. In Section II, we review the definition of Gaussian Boson sampling. We introduce the standard Glauber dynamics for sampling matchings in Section III, and then propose our double-loop Glauber dynamics for simulating GBS on graphs with provable guarantee in Section IV. We present all experimental results in Section V.
II Gaussian Boson Sampling for Graph Problems
@@
$\operatorname{Pr}(\bar{n})=\left|\operatorname{Perm}\left(T_{S}\right)\right|^{2}=\left|\sum_{\sigma\in\mathcal{S}_{N}}\prod_{i=1}^{N}T_{S_{i,\sigma(i)}}\right|^{2},$
where $\mathcal{S}_{N}$ is the set of all permutations on $[N]:=\{1,2,\ldots,N\}$, and $T_{S}$ is a matrix composed of the intersecting elements of the columns and the rows of $T$ determined by the input positions and output $\bar{n}$, respectively.
@@
Given an arbitrary undirected graph with potentially complex-valued adjacency matrix $\Delta$, we aim to engineer $B=c\Delta$ with an appropriate rescaling parameter $c$. It is possible to find such a $T$ by the Takagi-Autonne decomposition [20] when $0<c<1/({\max_{j}|\lambda_{j}|})$, where $\{\lambda_{j}\}$ is the eigenvalue set of $A$. Subsequently, the sampling matrix becomes $A=c\Delta\oplus c\Delta^{*}$.
@@
$(\sigma+\mathbb{I}_{2M}/2)^{-1}=\mathbb{I}_{2M}-\left(\begin{array}[]{ll}0_{M}&\mathbb{I}_{M}\\ \mathbb{I}_{M}&0_{M}\end{array}\right)A.$
@@
$\operatorname{Pr}_{A}(S)$ $=\sqrt{\det[(\sigma+\mathbb{I}_{2M}/2)^{-1}]}\operatorname{Haf}[(c\Delta\oplus c\Delta^{*})_{S}]$
$=\sqrt{\det\left(\begin{array}[]{cc}\mathbb{I}_{M}&-c\Delta^{*}\\ -c\Delta&\mathbb{I}_{M}\end{array}\right)}\operatorname{Haf}[c(\Delta_{S}\oplus\Delta_{S}^{\top})]$
$=\prod_{j=1}^{2M}\sqrt{1-c^{2}\lambda_{j}^{2}}c^{2|S|}\operatorname{Haf}^{2}(\Delta_{S}).$
In all, the task of simulating Gaussian Boson sampling on a graph is equivalent to developing algorithms that sample a subgraph with probability proportional to
@@
We remark that as a special case of the Hafnian, the permanent of a matrix with non-negative entries admits probabilistic polynomial-time approximation, in contrast to the #P-hardness for general complex matrices [19, 21, 22]. This suggests that the applications of GBS for max-Hafnian and densest $k$-subgraph of nonnegative-weight graphs may be simpler than the complex-valued version, and efficient classical algorithms for simulating GBS on such instances are worth investigation.
III Glauber Dynamics for Matching
Our algorithm is built upon the Glauber dynamics, a well-established Markov chain Monte Carlo method for sampling. In particular, the Glauber dynamics that samples across the space of all matchings of a graph is known as the monomer-dimer model. Given a graph $G=(V,E)$ and a fugacity parameter $\lambda>0$, we denote $\mathcal{M}$ to be the collection of all the matchings of $G$. We define the Gibbs distribution $\mu$ for the monomer-dimer model as $\mu(X)=w(X)/Z$ for $\forall X\in\mathcal{M}$, where the weight $w(X)=\lambda^{|X|}$ and $Z$ is a normalizing factor known as the partition function  [^1] This partition function is known as the matching polynomial, which has rich study in MCMC literature [25]..
@@ III Glauber Dynamics for Matching
1 Initialize $X_{0}$ as an arbitrary matching in $G$;
@@ 2 Initialize $t\leftarrow 0$;
3
4while $t<T$ do
5Choose a uniformly random edge $e$ from $E$;
@@
7Set $X_{t+1}=X_{t}\cup\{e\}$ with probability $\frac{\lambda}{1+\lambda}$ and otherwise set $X_{t+1}=X_{t}$;
8
9else if $e$ is in $X_{t}$ then
10Set $X_{t+1}=X_{t}\backslash\{e\}$ with probability $\frac{1}{1+\lambda}$ and otherwise set $X_{t+1}=X_{t}$;
@@ 12   Set $X_{t+1}=X_{t}$;
$\Pr[X\to X\cup\{e\}]=\frac{\lambda}{1+\lambda}\frac{1}{|E|},$
$\Pr[X\cup\{e\}\to X]=\frac{1}{1+\lambda}\frac{1}{|E|},$
@@
$\|P_{t}(X_{0},\cdot)-\mu\|_{\operatorname{TV}}:=\frac{1}{2}\sum_{X}|P_{t}(X_{0},X)-\mu(X)|.$
@@
$t_{\operatorname{mix}}(\epsilon)$ $:=\min\left\{t:\max_{X_{0}\in\mathcal{M}}\|P_{t}(X_{0},\cdot)-\mu\|_{\operatorname{TV}}\leq\epsilon\right\},$
$t_{\operatorname{mix}}$ $:=t_{\operatorname{mix}}({1}/{4}).$
@@
For a general graph $G$, the mixing time of the Glauber dynamics for the monomer-dimer model on $G$ with fugacity $\lambda>0$ is $O(n^{2}m\log n)$.
@@
$\|P_{t}^{\prime}(X_{0},\cdot)-\nu\|_{\operatorname{TV}}:=\frac{1}{2}\sum_{S}|P^{\prime}_{t}(X_{0},S)-\nu(S)|.$
@@
$t_{\operatorname{mix}}^{S}(\epsilon)$ $:=\min\left\{t:\max_{X_{0}\in\mathcal{M}}\|P^{\prime}_{t}(X_{0},\cdot)-\nu\|_{\operatorname{TV}}\leq\epsilon\right\},$
$t_{\operatorname{mix}}^{S}$ $:=t_{\operatorname{mix}}^{S}({1}/{4}).$
@@
$\|P^{\prime}_{t}(X_{0},\cdot)-\nu\|_{\operatorname{TV}}$
$=$ $\frac{1}{2}\sum_{S}|P^{\prime}_{t}(X_{0},S)-\nu(S)|$
$\leq$ $\frac{1}{2}\sum_{S}\sum_{X\text{ is a perfect matching of }S}|P_{t}(X_{0},X)-\mu(X)|$
$=$ $\|P_{t}(X_{0},\cdot)-\mu\|_{\operatorname{TV}},$
@@
IV Double-loop Glauber Dynamics
Inspired by the standard Glauber dynamics, we develop several enhanced algorithms that achieve classical simulation of GBS on graphs, i.e., sampling of a vertex set $S$ with distribution $\Pr(S)\propto c^{2|S|}\operatorname{Haf}^{2}(S)$.
A simple idea is rejection sampling, which is a basic technique applied to generate samples from a target distribution by sampling from a proposal distribution and accepting or rejecting the samples based on a certain criterion. For simulating GBS, our rejection sampling algorithm works as follows:
@@ IV Double-loop Glauber Dynamics
$\Pr[S\text{ is accepted}]$ $=\Pr[S_{1}=S]\cdot\Pr[S_{2}=S]$
$\propto({c}^{|S|}\operatorname{Haf}(S))^{2}=c^{2|S|}\operatorname{Haf}^{2}(S).$
@@
In this paper, we propose a novel double-loop Glauber dynamics that directly samples from the distribution $\Pr[S]\propto c^{2|S|}\operatorname{Haf}^{2}(S)$, where $S$ is a vertex set of a graph $G$ and $c$ is a constant. Note that it is equivalent to realizing the sampling of matchings according to the distribution $\Pr[X]\propto(c^{2})^{2|X|}\operatorname{Haf}(G_{X})$ , where $G_{X}$ denotes the subgraph induced by the vertex set of the matching $X$.
@@
2 Initialize $X_{0}$ as an arbitrary matching in $G$
3Initialize $t\leftarrow 0$, set $T=\tilde{O}(n^{6})$

### v2

# Efficient Classical Sampling from Gaussian Boson Sampling Distributions on Unweighted Graphs

@@ Abstract
[…]
@@ I Introduction
On the other hand, classical simulation algorithms for GBS have been actively explored. Quesada and Arrazola [13] introduced an exact classical algorithm for simulating GBS by sequentially sampling the number of photons in each mode conditioned on the previously sampled modes, which runs in polynomial space and exponential time. Oh et al. [14] developed a classical algorithm for Boson sampling based on dynamic programming, which exploits the limited connectivity of the linear-optical circuit to improve efficiency. Separately, Oh et al. [15] proposed a tensor-network-based classical algorithm to simulate large-scale GBS experiments with photon loss, requiring relatively modest computational resources. In the context of graph-theoretical applications of GBS, an important property is that the adjacency matrix encoded in the Gaussian Boson sampler is non-negative, which is believed to make the problem more tractable than in the general case. The aforementioned result [13] reduced the simulation of GBS with non-negative matrices to the problem of estimating Hafnians of non-negative matrices, which can be done efficiently for the adjacency matrix of certain strongly expanding graphs [16]. Oh et al. [17] also took advantage of this non-negativity property to design a quantum-inspired classical algorithm for finding dense subgraphs and their numerical results suggest that the advantage offered by a Gaussian Boson sampler is not significant. However, an open question remains whether a classical algorithm for GBS on general graphs with provable performance guarantees can achieve a computational cost comparable to that of a Gaussian Boson sampler.
@@
In this work, we adopt Markov chain Monte Carlo (MCMC) algorithms to sample from GBS distributions on unweighted graphs. MCMC is a standard class of sampling algorithms with well-established theoretical guarantees [18, 19]. Among MCMC methods, Glauber dynamics [20] is particularly widespread due to its simplicity and rigorous analytical foundations. Glauber dynamics generates samples from the matchings of an undirected and unweighted graph by iteratively adding or removing edges with biased transition probabilities. The resulting stationary distribution is proportional to the power of the number of edges in the matchings. For each matching sampled from Glauber dynamics, we consider its support vertex set as a sampled subset of vertices. The probability of sampling a given vertex set is further weighted by the number of perfect matchings within that set, which is also the Hafnian of the adjacency matrix of the graph.
@@
We propose a double-loop Glauber dynamics with a rigorous theoretical guarantee that its stationary distribution is identical to the sampling distribution of GBS on unweighted graphs. Specifically, unlike the standard single-loop Glauber dynamics, which yields a stationary distribution proportional to the Hafnian of subgraphs, the double-loop approach ensures a stationary distribution proportional to the square of the Hafnian, which coincides with the sampling distribution of GBS. Concretely, in the double-loop Glauber dynamics, when deciding whether to remove an edge, we run a secondary Markov chain to uniformly sample a perfect matching from the current subgraph. The removal decision is then based on whether this edge reappears in the newly sampled matching. Furthermore, for dense graphs, we prove that the double-loop Glauber dynamics has a polynomial mixing time, demonstrating its computational feasibility in these cases. The key to demonstrate the rapid mixing lies in the canonical path technique introduced in [18], which routed flows between every pair of matchings without creating particularly congested “pipes”, and bounded the mixing time by estimating the maximum congestion of possible transitions. More specifically, for complete graphs, we establish an enhanced canonical path framework that permits multiple paths between any two matchings. By constructing specially designed paths with favorable symmetry properties to distribute flows efficiently, we ultimately estimate maximum congestion through calculating total congestion of symmetrical transitions. For dense graphs, the maximum congestion is bounded by estimating the ratio between the congestion of dense graphs and complete graphs, thereby yielding a polynomial mixing time. This result is particularly significant because the dense graph represents the regime where classical methods are challenging. Sparse graphs often permit classical shortcuts that dense graphs lack. For instance, the Maximum Clique problem, a canonical task that GBS has been proposed to solve [21], is NP-hard in general. However, its classical complexity is greatly reduced on sparse graphs, where algorithms can exploit structural properties such as low degeneracy to find maximal cliques efficiently [22].
[…]
@@
The rest of the paper is organized as follows. In Section II.1, we review the definition of GBS. We introduce the standard Glauber dynamics for sampling matchings in Section II.2, and then propose our double-loop Glauber dynamics for sampling from GBS distributions on unweighted graphs with provable guarantee in Section II.3. We present all experimental results in Section II.4.
II Results
II.1 Gaussian Boson Sampling for graph problems
@@
$\operatorname{Pr}(\bar{n})=\left|\operatorname{Perm}\left(T_{S}\right)\right|^{2}=\left|\sum_{\sigma\in\mathcal{S}_{N}}\prod_{i=1}^{N}T_{S_{i,\sigma(i)}}\right|^{2},$ (1)
where $\mathcal{S}_{N}$ is the set of all permutations on $N]:=\{1,2,\ldots,N\}$, and $T_{S}$ is a matrix composed of the intersecting elements of the columns and the rows of $T$ determined by the input positions and output $\bar{n}$, respectively.
@@
Given an arbitrary undirected graph with potentially complex-valued symmetric adjacency matrix $\Delta$, we aim to engineer $B=c\Delta$ with an appropriate rescaling parameter $c$. It is possible to find such a $T$ by the Takagi-Autonne decomposition (see [24] and Section 4.4 of [25]) when $0<c<1/({\max_{j}|\lambda_{j}|})$, where $\{\lambda_{j}\}$ is the eigenvalue set of $A$. Subsequently, the sampling matrix becomes $A=c\Delta\oplus c\Delta^{*}$.
@@
$(\sigma+\mathbb{I}_{2M}/2)^{-1}=\mathbb{I}_{2M}-\left(\begin{array}[]{ll}0_{M}&\mathbb{I}_{M}\\ \mathbb{I}_{M}&0_{M}\end{array}\right)A.$ (7)
@@
$\operatorname{Pr}_{A}(S)$ $=\sqrt{\det[(\sigma+\mathbb{I}_{2M}/2)^{-1}]}\operatorname{Haf}[(c\Delta\oplus c\Delta^{*})_{S}]$ (8)
$=\sqrt{\det\left(\begin{array}[]{cc}\mathbb{I}_{M}&-c\Delta^{*}\\ -c\Delta&\mathbb{I}_{M}\end{array}\right)}\operatorname{Haf}(c\Delta_{S})\cdot\operatorname{Haf}(c\Delta_{S})^{*}$ (11)
$=\prod_{j=1}^{2M}\sqrt{1-c^{2}\lambda_{j}^{2}}c^{2|S|}|\operatorname{Haf}(\Delta_{S})|^{2}.$ (12)
In all, the task of sampling from GBS distributions on a graph with real-valued adjacency matrix is equivalent to developing algorithms that sample a subgraph with probability proportional to
@@
We remark that as a special case of the Hafnian, the permanent of a matrix with non-negative entries admits probabilistic polynomial-time approximation, in contrast to the #P-hardness for general complex matrices [23, 26, 27]. This suggests that the applications of GBS for max-Hafnian and densest $k$-subgraph of nonnegative-weight graphs may be simpler than the complex-valued version, and efficient classical sampling algorithms from GBS distributions on such instances are worth investigation.
II.2 Glauber dynamics for matching
Our algorithm is built upon the Glauber dynamics, a well-established Markov chain Monte Carlo method for sampling. In particular, the Glauber dynamics that samples across the space of all matchings of a graph is known as the monomer-dimer model. Given a graph $G=(V,E)$ and a fugacity parameter $\lambda>0$, we denote $\mathcal{M}$ to be the collection of all the matchings of $G$. We define the Gibbs distribution $\mu$ for the monomer-dimer model as $\mu(X)=w(X)/Z$ for $\forall X\in\mathcal{M}$, where the weight $w(X)=\lambda^{|X|}$ and $Z$ is a normalizing factor known as the partition function. (This partition function is known as the matching polynomial, which has rich study in MCMC literature [28]. The matching polynomial is closely related to the loop hafnian, a variant of the Hafnian that can be used to count all matchings in a graph [29, 30].)
@@ II.2 Glauber dynamics for matching
1
21em Initialize $X_{0}$ as an arbitrary matching in $G$;
@@ 3 Initialize $t\leftarrow 0$;
4
5while $t<T$ do
6Choose a uniformly random edge $e$ from $E$;
@@
8Set $X_{t+1}=X_{t}\cup\{e\}$ with probability $\frac{\lambda}{1+\lambda}$ and otherwise set $X_{t+1}=X_{t}$;
9
10else if $e$ is in $X_{t}$ then
[…]
@@ 13   Set $X_{t+1}=X_{t}$;
$\Pr[X\to X\cup\{e\}]=\frac{\lambda}{1+\lambda}\frac{1}{|E|},$ (14)
$\Pr[X\cup\{e\}\to X]=\frac{1}{1+\lambda}\frac{1}{|E|},$ (15)
@@
$\|P_{t}(X_{0},\cdot)-\mu\|_{\operatorname{TV}}:=\frac{1}{2}\sum_{X}|P_{t}(X_{0},X)-\mu(X)|.$ (17)
@@
$t_{\operatorname{mix}}(\epsilon)$ $:=\min\left\{t:\max_{X_{0}\in\mathcal{M}}\|P_{t}(X_{0},\cdot)-\mu\|_{\operatorname{TV}}\leq\epsilon\right\},$ (18)
$t_{\operatorname{mix}}$ $:=t_{\operatorname{mix}}({1}/{4}).$ (19)
@@
For a general graph $G$ with $n$ vertices and $m$ edges, the mixing time of the Glauber dynamics for the monomer-dimer model on $G$ with fugacity $\lambda>0$ is $O(n^{2}m\log n)$.
@@
$\|P_{t}^{\prime}(X_{0},\cdot)-\nu\|_{\operatorname{TV}}:=\frac{1}{2}\sum_{S}|P^{\prime}_{t}(X_{0},S)-\nu(S)|.$ (22)
@@
$t_{\operatorname{mix}}^{S}(\epsilon)$ $:=\min\left\{t:\max_{X_{0}\in\mathcal{M}}\|P^{\prime}_{t}(X_{0},\cdot)-\nu\|_{\operatorname{TV}}\leq\epsilon\right\},$ (23)
$t_{\operatorname{mix}}^{S}$ $:=t_{\operatorname{mix}}^{S}({1}/{4}).$ (24)
@@
$\|P^{\prime}_{t}(X_{0},\cdot)-\nu\|_{\operatorname{TV}}$ (25)
$=$ $\frac{1}{2}\sum_{S}|P^{\prime}_{t}(X_{0},S)-\nu(S)|$ (26)
$\leq$ $\frac{1}{2}\sum_{S}\sum_{X\text{ is a perfect matching of }S}|P_{t}(X_{0},X)-\mu(X)|$ (27)
$=$ $\|P_{t}(X_{0},\cdot)-\mu\|_{\operatorname{TV}},$ (28)
@@
II.3 Double-loop Glauber dynamics
Inspired by the standard Glauber dynamics, we develop several enhanced algorithms that achieve classical sampling from GBS distributions on unweighted graphs, i.e., sampling of a vertex set $S$ with distribution $\Pr(S)\propto c^{2|S|}\operatorname{Haf}^{2}(S)$.
A simple idea is rejection sampling, which is a basic technique applied to generate samples from a target distribution by sampling from a proposal distribution and accepting or rejecting the samples based on a certain criterion. For sampling from GBS distributions, our rejection sampling algorithm works as follows:
@@ II.3 Double-loop Glauber dynamics
$\Pr[S\text{ is accepted}]$ $=\Pr[S_{1}=S]\cdot\Pr[S_{2}=S]$ (29)
$\propto({c}^{|S|}\operatorname{Haf}(S))^{2}=c^{2|S|}\operatorname{Haf}^{2}(S).$ (30)
@@
In this work, we propose a novel double-loop Glauber dynamics that directly samples from the distribution $\Pr[S]\propto c^{2|S|}\operatorname{Haf}^{2}(S)$, where $S$ is a vertex set of a graph $G$ and $c$ is a constant. Note that it is equivalent to realizing the sampling of matchings according to the distribution $\Pr[X]\propto(c^{2})^{2|X|}\operatorname{Haf}(G_{X})$ , where $G_{X}$ denotes the subgraph induced by the vertex set of the matching $X$.
@@
2 Initialize $X_{0}$ as an arbitrary matching in $G$;
3Initialize $t\leftarrow 0$, set $T=\tilde{O}(n^{6})$;